# Flexible inference in heterogeneous and attributed multilayer networks

### v1

@@ Results
We assess the performance of PIHAM by evaluating its predictive capabilities through a $5$-fold cross-validation routine. In this approach, the dataset is randomly divided into five equal-sized groups (folds), and the model is trained on four of them (the training set), which include $80\%$ of the triples $(i,j,\ell)$ and $80\%$ of the entries of each attribute vector, to learn its parameters. The performance of the model is then evaluated on the remaining fold (the test set). This process is repeated five times, each time with a different fold as the test set, resulting in five trials per iteration. For performance metrics, we use different measures depending on the type of information being evaluated. For binary interactions, we use the Area Under the Curve (AUC), which represents the probability that a randomly selected edge has a higher expected value than a randomly selected non-existing edge. The AUC ranges from 00 to $1$, with $0.5$ representing the random baseline. For nonnegative discrete data, we use the Maximum Absolute Error (MAE), and for real values, we use the Root Mean Squared Error (RMSE). In both cases, lower values indicate better performance. Additionally, for categorical attribute predictions, we use accuracy, which ranges from 00 to $1$, with $1$ indicating perfect recovery.
@@
To determine the number of communities $K$, we employ a $5$-fold cross-validation procedure for $K\in[1,10]$ and select the value that exhibits the optimal performance. Similar to the synthetic experiments, for a given $K$, we train the model on four folds (training set), which include $80\%$ of the triples $(i,j,\ell)$ and $80\%$ of the entries of each attribute vector, to learn its parameters. The model’s performance is then evaluated on the remaining fold (the test set). This process is repeated five times, each with a different fold as the test set, resulting in five trials per iteration. We evaluate prediction performance in the first six binary layers using the Area Under the Curve (AUC), representing the probability that a randomly selected edge has a higher expected value than a randomly selected non-existing edge. The AUC ranges from 00 to $1$, with $0.5$ indicating the random baseline. For the seventh layer containing real values, we use the Root Mean Squared Error (RMSE), where lower values indicate better performance. Additionally, we assess prediction performance for the attributes using accuracy for caste and religion, which ranges from 00 to $1$ (with $1$ indicating perfect recovery), and the Maximum Absolute Error (MAE) for years of education, with lower values indicate better performance. The results are displayed in Table S3. In our experiments, we set $K=6$ as it achieves the best performance across most prediction metrics. Note that, summarizing and evaluating results in a heterogeneous setting using a single metric is challenging, as discussed in the section Validation on heterogeneous data. of the main text. The results in Table S3 further demonstrate that PIHAM achieves robust outcomes with the chosen fixed value of $K$.
@@
Taken together, these findings suggest that PIHAM utilizes all the input information to infer partitions that effectively integrate all of them in a meaningful manner. In addition, the inferred affinity matrices illustrate how different layers can exhibit different community structures, a diversity that can be captured by our model.

### v2

@@ Results
We assess the performance of PIHAM by evaluating its predictive capabilities through a $5$-fold cross-validation routine. In this approach, the dataset is randomly divided into five equal-sized groups (folds), and the model is trained on four of them (the training set), which include $80\%$ of the triples $(i,j,\ell)$ and $80\%$ of the entries of each attribute vector, to learn its parameters. The performance of the model is then evaluated on the remaining fold (the test set). This process is repeated five times, each time with a different fold as the test set, resulting in five trials per iteration. For performance metrics, we use different measures depending on the type of information being evaluated. For binary interactions, we use AUC, which ranges from 00 to $1$, with $0.5$ representing the random baseline. For nonnegative discrete data, we use the Maximum Absolute Error (MAE), and for real values, we use the Root Mean Squared Error (RMSE). In both cases, lower values indicate better performance. Additionally, for categorical attribute predictions, we use accuracy, which ranges from 00 to $1$, with $1$ indicating perfect recovery.
@@
To determine the number of communities $K$, we employ a $5$-fold cross-validation procedure for $K\in[1,10]$ and select the value that exhibits the optimal performance. Similar to the synthetic experiments, for a given $K$, we train the model on four folds (training set), which include $80\%$ of the triples $(i,j,\ell)$ and $80\%$ of the entries of each attribute vector, to learn its parameters. The model’s performance is then evaluated on the remaining fold (the test set). This process is repeated five times, each with a different fold as the test set, resulting in five trials per iteration. We evaluate prediction performance in the first six binary layers using the AUC, ranging from 00 to $1$, with $0.5$ indicating the random baseline. For the seventh layer containing real values, we use the RMSE, where lower values indicate better performance. Additionally, we assess prediction performance for the attributes using accuracy for caste and religion, which ranges from 00 to $1$ (with $1$ indicating perfect recovery), and the MAE for years of education, with lower values indicate better performance. The results are displayed in Table S3. In our experiments, we set $K=6$ as it achieves the best performance across most prediction metrics. Note that, summarizing and evaluating results in a heterogeneous setting using a single metric is challenging, as discussed in the section Validation on heterogeneous data. of the main text. The results in Table S3 also demonstrate that PIHAM achieves robust outcomes with the chosen fixed value of $K$.
We further assess whether the model is a good fit to the data by conducting a posterior–predictive test [41, 42], where we compare the input data $(\boldsymbol{A},\boldsymbol{X})$ to synthetic data $(\tilde{\boldsymbol{A}},\tilde{\boldsymbol{X}})$ generated by the fitted model. A well-fitted model should produce synthetic data that closely match the original input. The posterior-predictive distribution is defined as
[…]
and samples from this distribution are obtained by first drawing $\boldsymbol{\Theta}$ from the posterior distributions $\mathcal{N}(\boldsymbol{\theta};\hat{\boldsymbol{\mu}}^{\boldsymbol{\theta}},\hat{\boldsymbol{\Sigma}}^{\boldsymbol{\theta}})$, and then using these parameters to create new synthetic data. We apply the likelihood described in the previous subsection, fixing the standard deviation to $0.1$ for the Gaussian layer, and generate $500$ independent synthetic datasets. To evaluate performance, we test whether two samples from the posterior-predictive distribution are generally more, equally, or less distant from each other than a sample from the posterior-predictive distribution compared to the input data [41]. We measure the distance using different metrics depending on the data type: log-loss for the six binary layers; RMSE for the seventh layer with real values, $(1-\text{accuracy})$ for categorical attributes, and MAE for the attribute representing years of education. The results in Fig. S7 show that the discrepancies between synthetic data samples consistently exceed those between the observed data and synthetic samples, indicating that PIHAM  provides a good fit for the data.
@@
Fig. S7: Goodness of fit test for the “Alakāpuram” social support network. We generate $500$ independent synthetic datasets using the fitted model and compare them to both the input data and other synthetic samples. Each dot represents one posterior-predictive sample, showing the distance between this sample and the input data on the horizontal axis, and a randomly selected synthetic sample on the vertical axis. Performance is evaluated using different metrics $D$ depending on the data type: log-loss for binary interactions (A), RMSE for real values (B), $(1-\text{accuracy})$ for categorical attributes (D, E), and MAE for nonnegative discrete values (F). The differences between synthetic data samples consistently exceed those between the observed data and synthetic samples, as all points lie above the diagonal, indicating that PIHAM provides a good fit for the data.